OpenStudy (anonymous):

what is the slope of the line described the equation y-7=8(x-14) is the answer -7 ?

OpenStudy (abb0t):

It's not -7 NOTE: m = slope, where: y =mx + b and similarly \[(y-y_1)=m(x-x_1)\]

OpenStudy (anonymous):

8!

OpenStudy (abb0t):

Correct! Good job.